OpenStudy (anonymous):

7x=5y Find the slope and the y intercepts of each line whose equation is given?

OpenStudy (anonymous):

You want to put it in Slope-Intercept Form, which is y=mx+b. m=Slope b=y-intercept

OpenStudy (anonymous):

ok y intercept is 0?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

so how do i get from 7x=5y to Y=(7/5)x+0

OpenStudy (anonymous):

You're just isolating the y by itself... so you divide each side by 5. That leaves you 7/5x=y

OpenStudy (anonymous):

how do i lose the x?

OpenStudy (anonymous):

You don't. The x is part of the ordered pair (x,y) for the equation.

OpenStudy (anonymous):

Are you still looking for the slope?

OpenStudy (anonymous):

no i just know that 7/5 is the slope and 0 is the intercept.. didnt know if the x is 7/5 and y is o?

OpenStudy (anonymous):

No. It's 7/5 times x equals y. In the full Slope-Intercept Form, it's 7/5x+0=y Where 7/5 is the slope, and 0 is the y-intercept.

OpenStudy (anonymous):

ok so sorry i dont get this stuff. if 7/5 is times x, where did x go? is x =1?

OpenStudy (anonymous):

x is whatever. the x and y in the equation is any ordered pair. If you're only looking for the y-intercept and the slope, the x doesn't matter. As long as you realize that the slope is coefficient of x.

OpenStudy (anonymous):

oh ok. got it.. thanks.. can u help with this one too? this is my last week in principles in algebra and i have to pass to go on to next class. I promise, I will not get a job requiring this stuff!! lol here is problem....Write an equation of the line containing the specified point and perpendicular to the indicated line (4, −2), x = 1

OpenStudy (anonymous):

Okay. Sorry, this is taking me a minute. lol

OpenStudy (anonymous):

oh ok.. lol i understand and this isnt even my final! lol just two questions..

OpenStudy (anonymous):

For lines to be perpendicular, the slope of line one times the slope of line two must equal -1.

OpenStudy (anonymous):

ok? lol well i have to go attend to the kiddos..lol i will have to come back tomorrow for more help.. if you are here i will holla at ya! thanks so much for your help so far!

OpenStudy (anonymous):

No problem. Glad I could help a little!